# Machine Learning-Based Detection of Pump-and-Dump Schemes in Real-Time

### v1

@@ V-A2 Price Spikes
Figure 5: Relative Price Changes and Price Spike Metrics During KuCoin Pump Events.
Figure 6: Relative Price Changes and Price Spike Metrics During Poloniex Pump Events.
On average, the time to reach the maximum price spike was calculated at 1.49 minutes for KuCoin and just 0.12 minutes for Poloniex. The magnitude of the price spikes, measured relative to a baseline (the midpoint between the bid and ask prices ten minutes prior to the pump), was notably higher on Poloniex (10.85) compared to KuCoin (4.57). These results suggest that pumps on Poloniex not only reach their peak more quickly but also result in larger relative price spikes. This difference aligns with the observations from supplementary Figures 13 and 13 that show lower trading volumes and liquidity on Poloniex. Limited market depth and lower liquidity make it easier to induce extreme price fluctuations, leading to shorter-lived but sharper price spikes during pump events.
V-A3 Order Size
@@
On KuCoin, selected coins exhibit varying trends in order size during pump events. As shown in Table IV, KOL has the largest average order size before the pump, suggesting early trader positioning in anticipation of the event. Conversely, KPOL displays a significant increase in average order size during the pump, with an increase ratio of 1.30, indicating heightened interest and aggressive buying activity. IZI, pumped twice, shows a consistent decline in order size during the pump, likely due to profit-taking or cautious trading, as traders place smaller orders to mitigate risks in volatile conditions.
TABLE IV: Average Order Sizes for Selected Coins on KuCoin During the 7 Days Before the Pump and on the Pump Day
@@
On Poloniex, market behavior during pump events is notably more volatile. Table V highlights varied trends, with some coins displaying large order sizes before the pump, indicative of early positioning, while others experience sharp increases during the pump, reflecting sudden trader interest. For example, AMC shows inconsistent behavior, with stable order sizes in one instance and a dramatic increase in another, suggesting unpredictable trader dynamics. DMT and GFT demonstrate cautious trading, with declines in order size during the pump. Many coins on Poloniex experience periods of inactivity, signaling market hesitation or uncertainty leading up to the pump.
TABLE V: Average Order Sizes for Selected Coins on Poloniex During the 7 Days Before the Pump and on the Pump Day
@@
Supplementary Figures 13 and 13 in Appendix visualize the differences in average order sizes across both exchanges. KuCoin exhibits more predictable patterns, with traders often positioning themselves ahead of pump events, as seen with KOL. In contrast, Poloniex shows greater variability, with some coins like COLLAB maintaining consistent order sizes, while others, such as AMC and GFT, demonstrate dramatic fluctuations during the pump.
V-B Filter Models to Reduce the Candidate Pool
Cryptocurrency exchanges often list thousands of tokens, creating significant scalability challenges for real-time tracking. For instance, MEXC lists 2,738 tokens [33]. To manage this complexity, our pipeline calculates and stores only order book and trade features in real-time, avoiding the need to store raw data. However, to further reduce the number of tokens tracked, we developed filter models to eliminate coins that lack the attributes commonly associated with pumped tokens.
@@ VI-A Interpretation of Empirical Results
Future iterations could benefit from integrating on-chain transaction data, as P&D organizers often transfer liquidity from other exchanges to the pump’s target platform. Monitoring large or suspicious deposits into the target exchange could serve as an early warning signal. If on-chain data is unavailable, expanding the analysis to include multiple exchanges could improve detection capabilities, albeit with increased computational demands. Additionally, incorporating broader market indicators, such as social media sentiment and external events, could enhance early detection accuracy. Addressing data sparsity through a more robust comparison pool is another priority. Techniques such as jump analysis [34] or fine-tuning time series transformer models [35] could further refine anomaly detection. Leveraging pretrained time series models with labeled datasets (see Appendix E-A) offers another promising avenue for improvement.
VI-B Susceptibility of Token Standards for P&D Schemes
This section presents how various token standards influence their susceptibility to P&D schemes. By analyzing the implementation of non-fungible token standards such as ERC-721 [24] with fungible token standards like ERC-20 [21] and Runes [11], we assessed how their properties allow or prevent from being exploited.
VI-B1 Non-Fungible Token Standards
Non-fungible token (NFT) standards, such as ERC-721 tokens and Inscriptions [25], are generally unsuitable for P&D schemes. NFTs are unique and often linked to a specific digital or physical asset, making them incompatible with the mass trading needed for P&D events. Unlike fungible tokens, where multiple parties can hold large quantities simultaneously, only one entity can own an NFT at any given time.
@@ VI-B1 Non-Fungible Token Standards
Additionally, certain operational barriers exist for trading NFTs like Inscriptions, which require the ”ord” client and a fully synchronized Bitcoin node [36]. Supporting this analysis, our historical dataset revealed no instances of NFTs being targeted in P&D schemes. Telegram channels and marketplaces for NFTs, such as OpenSea and Magic Eden, lack any signs of coordinated P&D activity. Fragmented ownership of NFTs, which enables shared liquidity, might create conditions more conducive to P&D schemes. While not observed in our data, other manipulation tactics, such as wash trading—where the same entity repeatedly buys and sells an NFT to inflate its price—are more likely in the NFT space.
VI-B2 Fungible Token Standards
In contrast to NFTs, fungible tokens are inherently more susceptible to P&D schemes due to their interchangeable nature. Standards such as ERC-20, BEP-20 [37], and experimental types like BRC-20 [38] or Runes enable mass trading, making them ideal for orchestrators seeking to manipulate prices. These tokens allow multiple traders to hold large amounts simultaneously, facilitating coordinated buying during the pump phase and subsequent dumping for profit. ERC-20-like tokens, in particular, are widely used due to their ease of listing on exchanges and large market volumes. Their relatively low entry barriers for traders and higher anonymity compared to NFTs make them prime targets for manipulation. Anonymity, afforded by high token supplies, conceals orchestrators’ actions and ensures smoother coordination.
All target coins in observed P&D schemes in this study adhered to ERC-20-like standards. However, no pumps were observed involving listed BRC-20 tokens or Runes. This can be attributed to their limited trading availability on centralized exchanges [39], their experimental nature, and the lower adoption rates. Additionally, BRC-20 tokens face inefficiencies, including slower transaction speeds and higher fees, especially under network congestion [10]. These barriers reduce their attractiveness for orchestrators seeking rapid and large-scale price manipulation.
@@ Appendix A Appendix: Natural Language Processing
A-A OpenAI Settings and Prompts
@@ Appendix B Computed Metrics
B-A Order Book and Trade Data Metrics
Table IX defines the key metrics derived from order book and trade data, including their formulas and a brief explanation.
TABLE IX: Computed Metrics for Orderbook and Trade Data
@@ C-A XGBoost Model
C-A1 Feature Selection
From the initial 27 time series features, we calculated multiple features per time series using the Python library TS Fresh, as XGBoost cannot work with data in time series format. We applied recursive feature elimination with a 5-fold cross-validation, using the weighted F1 score as the evaluation metric. Table X lists the best features selected for each XGBoost model described in Section V-B.
TABLE X: Features and Descriptions for XGBoost Models 0 to 3
@@ C-A1 Feature Selection
C-A2 Hyperparameter Tuning
After feature selection, we trained the XGBoost classifier on various feature combinations. Hyperparameters, including max_depth, min_child_weight, and scale_pos_weight, were fine-tuned using grid search with 3-fold cross-validation. Interestingly, the default parameters of the XGBClassifier in the Python xgboost library performed better than the grid-searched parameters.
C-A3 Performance Evaluation
Figure 14: Confusion Matrices of XGBoost Models for Different Features
We evaluated performance using metrics like precision, recall, F1 score, and Precision-Recall AUC (Figures 15, 16, and 17). Table X highlights the reliance of all models on market data, such as maximum and minimum closing prices and market cap metrics. While social data (e.g., sentiment scores and user interactions) and geometric time series features (e.g., spectral entropy) marginally improved models 2 and 3, traditional market data remained the primary predictive factor.
Figure 15: Precision-Recall Curve for Models (Model 0: price_minimal, Model 1: price_optimal, Model 2: price_social, Model 3: price_social_geometric)
Figure 16: Weighted F1 Score vs. Threshold for Models
Figure 17: Precision-Recall AUC vs. Threshold for Models
C-B Unsupervised Clustering Experiments
@@ C-B1 Data Preprocessing
C-B2 DBSCAN (Density-Based Spatial Clustering of Applications with Noise)
DBSCAN groups points based on density while classifying isolated points as noise. Unlike other methods, it does not require specifying the number of clusters. Instead, it uses two parameters: min_samples, the minimum points to form a dense region, and eps, the radius of each neighborhood [40].
C-B3 K-means Clustering
K-means is widely used for dividing datasets into $k$ clusters. The algorithm iteratively updates cluster centers (centroids) and reassigns points until convergence [41]. In our study, K-means grouped coins based on historical data patterns. Multiple values of $k$ were tested to identify performance tiers among coins.
C-B4 Agglomerative Clustering
Agglomerative clustering is a hierarchical method where each data point starts as its own cluster, merging iteratively based on a chosen linkage criterion [42]. This method revealed multi-level hierarchies, helping visualize how clusters evolve and merge as distance thresholds increase.
C-B5 Gaussian Mixture Model (GMM)
GMM is a probabilistic approach that assumes data originates from multiple Gaussian distributions. Parameters for each cluster are estimated using the Expectation-Maximization (EM) algorithm [43]. GMM is particularly useful for handling clusters with varying shapes and sizes.
TABLE XI: Clustering Algorithm Performance Metrics
@@ C-B5 Gaussian Mixture Model (GMM)
Figure 18: Comparison of Different Clustering Methods

### v2

@@ 5.1.2. Price Spikes
Figure 5. Relative Price Changes and Price Spike Metrics During KuCoin Pump Events.
Figure 6. Relative Price Changes and Price Spike Metrics During Poloniex Pump Events.
On average, the time to reach the maximum price spike was calculated at 1.49 minutes for KuCoin and just 0.12 minutes for Poloniex. The magnitude of the price spikes, measured relative to a baseline (the midpoint between the bid and ask prices ten minutes prior to the pump), was notably higher on Poloniex (10.85) compared to KuCoin (4.57). These results suggest that pumps on Poloniex not only reach their peak more quickly but also result in larger relative price spikes. This difference aligns with the observations from supplementary Figures 18 and 18 that show lower trading volumes and liquidity on Poloniex. Limited market depth and lower liquidity make it easier to induce extreme price fluctuations, leading to shorter-lived but sharper price spikes during pump events.
5.1.3. Order Size
@@
On KuCoin, selected coins exhibit varying trends in order size during pump events. As shown in Table 5, KOL has the largest average order size before the pump, suggesting early trader positioning in anticipation of the event. Conversely, KPOL displays a significant increase in average order size during the pump, with an increase ratio of 1.30, indicating heightened interest and aggressive buying activity. IZI, pumped twice, shows a consistent decline in order size during the pump, likely due to profit-taking or cautious trading, as traders place smaller orders to mitigate risks in volatile conditions.
Table 5. Average Order Sizes for Selected Coins on KuCoin During the 7 Days Before the Pump and on the Pump Day
@@
On Poloniex, market behavior during pump events is notably more volatile. Table 6 highlights varied trends, with some coins displaying large order sizes before the pump, indicative of early positioning, while others experience sharp increases during the pump, reflecting sudden trader interest. For example, AMC shows inconsistent behavior, with stable order sizes in one instance and a dramatic increase in another, suggesting unpredictable trader dynamics. DMT and GFT demonstrate cautious trading, with declines in order size during the pump. Many coins on Poloniex experience periods of inactivity, signaling market hesitation or uncertainty leading up to the pump.
Table 6. Average Order Sizes for Selected Coins on Poloniex During the 7 Days Before the Pump and on the Pump Day
@@
Supplementary Figures 18 and 18 in Appendix visualize the differences in average order sizes across both exchanges. KuCoin exhibits more predictable patterns, with traders often positioning themselves ahead of pump events, as seen with KOL. In contrast, Poloniex shows greater variability, with some coins like COLLAB maintaining consistent order sizes, while others, such as AMC and GFT, demonstrate dramatic fluctuations during the pump.
5.2. Filter Models to Reduce the Candidate Pool
Cryptocurrency exchanges often list thousands of tokens, creating significant scalability challenges for real-time tracking. For instance, MEXC lists 2,738 tokens (Coinranking, 2024). To manage this complexity, our pipeline calculates and stores only order book and trade features in real-time, avoiding the need to store raw data. However, to further reduce the number of tokens tracked, we developed filter models to eliminate coins that lack the attributes commonly associated with pumped tokens.
@@ 6. Discussion
Future iterations could benefit from integrating on-chain transaction data, as P&D organizers often transfer liquidity from other exchanges to the pump’s target platform. Monitoring large or suspicious deposits into the target exchange could serve as an early warning signal. If on-chain data is unavailable, expanding the analysis to include multiple exchanges could improve detection capabilities, albeit with increased computational demands. Additionally, incorporating broader market indicators, such as social media sentiment and external events, could enhance early detection accuracy.
Addressing data sparsity through a more robust comparison pool is another priority. Techniques such as jump analysis (Lee and Mykland, 2008) or fine-tuning time series transformer models (Rasul et al., 2024) could further refine anomaly detection. Leveraging pretrained time series models with labeled datasets offers another promising avenue for improvement.
A recent advancement in time series forecasting is the development of the Lag LLAMA model (Rasul et al., 2024), which is based on the transformer architecture. This framework underlies sophisticated large language models such as ChatGPT and Claude. Transformer models excel at capturing long-range dependencies in sequential data, making them particularly suitable for a variety of predictive tasks.
Lag LLAMA has demonstrated promising results in predicting financial data (Bahelka and de Weerd, 2024). However, its current application is restricted to univariate data, which limits its utility for studies such as ours that involve multivariate datasets. Future extensions of the Lag LLAMA model to support multivariate data would open the door to new opportunities. Exploring its performance in these more complex, noisy, and interdependent settings could yield valuable insights and further refine predictive capabilities.
Susceptibility of Token Standards for P&D Schemes
Non-Fungible Token Standards. Non-fungible token (NFT) standards, such as ERC-721 tokens and Inscriptions (Rodarmor, 2023a), are generally unsuitable for P&D schemes. NFTs are unique and often linked to a specific digital or physical asset, making them incompatible with the mass trading needed for P&D events. Unlike fungible tokens, where multiple parties can hold large quantities simultaneously, only one entity can own an NFT at any given time.
@@
Additionally, certain operational barriers exist for trading NFTs like Inscriptions, which require the ”ord” client and a fully synchronized Bitcoin node (Nervos Network, 2024). Supporting this analysis, our historical dataset revealed no instances of NFTs being targeted in P&D schemes. Telegram channels and marketplaces for NFTs, such as OpenSea and Magic Eden, lack any signs of coordinated P&D activity. Fragmented ownership of NFTs, which enables shared liquidity, might create conditions more conducive to P&D schemes. While not observed in our data, other manipulation tactics, such as wash trading—where the same entity repeatedly buys and sells an NFT to inflate its price—are more likely in the NFT space.
Fungible Token Standards. In contrast to NFTs, fungible tokens are inherently more susceptible to P&D schemes due to their interchangeable nature. Standards such as ERC-20, BEP-20 (Binance, nd), and experimental types like BRC-20 (Domo, 2023) or Runes enable mass trading, making them ideal for orchestrators seeking to manipulate prices. These tokens allow multiple traders to hold large amounts simultaneously, facilitating coordinated buying during the pump phase and subsequent dumping for profit. ERC-20-like tokens, in particular, are widely used due to their ease of listing on exchanges and large market volumes. Their relatively low entry barriers for traders and higher anonymity compared to NFTs make them prime targets for manipulation. Anonymity, afforded by high token supplies, conceals orchestrators’ actions and ensures smoother coordination.
All target coins in observed P&D schemes in this study adhered to ERC-20-like standards. However, no pumps were observed involving listed BRC-20 tokens or Runes. This can be attributed to their limited trading availability on centralized exchanges (Team, 2024), their experimental nature, and the lower adoption rates. Additionally, BRC-20 tokens face inefficiencies, including slower transaction speeds and higher fees, especially under network congestion (Li et al., 2024). These barriers reduce their attractiveness for orchestrators seeking rapid and large-scale price manipulation.
@@ Appendix A Appendix: Natural Language Processing
A.1. OpenAI Settings and Prompts
@@ Appendix B Computed Metrics
B.1. Order Book and Trade Data Metrics
Table 7 defines the key metrics derived from order book and trade data, including their formulas and a brief explanation.
Table 7. Computed Metrics for Orderbook and Trade Data
@@ C.1. XGBoost Model
C.1.1. Feature Selection
From the initial 27 time series features, we calculated multiple features per time series using the Python library TS Fresh, as XGBoost cannot work with data in time series format. We applied recursive feature elimination with a 5-fold cross-validation, using the weighted F1 score as the evaluation metric. Table 8 lists the best features selected for each XGBoost model described in Section 5.2.
Table 8. Features and Descriptions for XGBoost Models 0 to 3
@@ C.1.1. Feature Selection
C.1.2. Hyperparameter Tuning
After feature selection, we trained the XGBoost classifier on various feature combinations. Hyperparameters, including max_depthmin_child_weight,
and scale_pos_weight, were fine-tuned using grid search with 3-fold cross-validation. Interestingly, the default parameters of the XGBClassifier in the Python xgboost library performed better than the grid-searched parameters.
C.1.3. Performance Evaluation
Figure 7. Confusion Matrices of XGBoost Models for Different Features
We evaluated performance using metrics like precision, recall, F1 score, and Precision-Recall AUC (Figures 8, 9, and 10). Table 8 highlights the reliance of all models on market data, such as maximum and minimum closing prices and market cap metrics. While social data (e.g., sentiment scores and user interactions) and geometric time series features (e.g., spectral entropy) marginally improved models 2 and 3, traditional market data remained the primary predictive factor.
Figure 8. Precision-Recall Curve for Models (Model 0: price_minimal, Model 1: price_optimal, Model 2: price_social, Model 3: price_social_geometric)
Figure 9. Weighted F1 Score vs. Threshold for Models
Figure 10. Precision-Recall AUC vs. Threshold for Models
C.2. Unsupervised Clustering Experiments
@@ C.2.1. Data Preprocessing
C.2.2. DBSCAN (Density-Based Spatial Clustering of Applications with Noise)
DBSCAN groups points based on density while classifying isolated points as noise. Unlike other methods, it does not require specifying the number of clusters. Instead, it uses two parameters: min_samples, the minimum points to form a dense region, and eps, the radius of each neighborhood (Ester et al., 1996).
C.2.3. K-means Clustering
K-means is widely used for dividing datasets into $k$ clusters. The algorithm iteratively updates cluster centers (centroids) and reassigns points until convergence (MacQueen, 1967). In our study, K-means grouped coins based on historical data patterns. Multiple values of $k$ were tested to identify performance tiers among coins.
C.2.4. Agglomerative Clustering
Agglomerative clustering is a hierarchical method where each data point starts as its own cluster, merging iteratively based on a chosen linkage criterion (Murtagh and Contreras, 2012). This method revealed multi-level hierarchies, helping visualize how clusters evolve and merge as distance thresholds increase.
C.2.5. Gaussian Mixture Model (GMM)
GMM is a probabilistic approach that assumes data originates from multiple Gaussian distributions. Parameters for each cluster are estimated using the Expectation-Maximization (EM) algorithm (Reynolds, 2009). GMM is particularly useful for handling clusters with varying shapes and sizes.
Table 9. Clustering Algorithm Performance Metrics
@@ C.2.5. Gaussian Mixture Model (GMM)
Figure 11. Comparison of Different Clustering Methods